# Rational motives on pro-algebraic stacks

### v3

@@ 2.3 Six functor formalisms
Note that by design we have $\widetilde{{\mathcal{D}}}_{|\operatorname{Fun}(K,{\mathcal{C}}^{\operatorname{op}})}\simeq{\mathcal{D}}_{K}$. We will construct a functor $F\colon K\times\Delta^{1}\rightarrow\textup{Corr}({{\mathcal{C}}})_{E}$ that automatically yields a colimit preserving functor $f_{!}$ satisfying the necessary properties via $\widetilde{{\mathcal{D}}}$.
Let $\sigma\coloneqq(\sigma_{K},\sigma_{0,1})$ be an $n$-simplex in $K\times\Delta^{1}$. If $\sigma_{0,1}$ is a degeneration of $0\in\Delta^{1}$, then we define $F(\sigma)\coloneqq X(\sigma_{K})$ and if it is a degeneration of $1\in\Delta^{1}$, then we define $F(\sigma)\coloneqq Y(\sigma_{K})$.
Now assume that $\sigma_{0,1}$ is not a degeneration of a 00-simplex. Let $d$ denote the face map that selects the final vertex of an $n$-simplex. We will construct a correspondence $C_{X,Y}$ inside $\textup{Corr}({{\mathcal{C}}})_{E}$ as follows. As before, we let $f_{d(\sigma_{K})}$ be the morphism $f(d(\sigma_{K}),\langle 0,1\rangle)\colon X(d(\sigma_{K}))\rightarrow Y(d(\sigma_{K}))$ in ${\mathcal{C}}$. This defines a morphism $\alpha$ of correspondences
${X(d_{f}(\sigma_{K}))}$${X(d_{f}(\sigma_{K}))}$${Y(d_{f}(\sigma_{K})).}$$\scriptstyle{f_{d(\sigma_{K})}}$
We let $d^{c}$ denote the face map that forgets about the final vertex, i.e. $d^{c}$ is induced by the map $[n-1]\mapsto[n]$ inside $\Delta$ whose image omits the element $n$. Then $Y(d^{c}(\sigma_{K}))$ defines a $(n-1)$-simplex $\beta$ in $\textup{Corr}({{\mathcal{C}}})_{E}$ via the inclusion ${\mathcal{C}}^{\operatorname{op}}\hookrightarrow\textup{Corr}({{\mathcal{C}}})_{E}$. Informally, we may think of $\beta$ as the composition of $Y(d_{n-2,n-1}(\sigma_{K}))^{*}Y(d_{n-3,n-2}(\sigma_{K}))^{*}\dots Y(d_{0,1}(\sigma_{K}))^{*}$, where the $d_{i,j}$ denotes the edge connecting the $i$-th and $j$-th vertex inside $d^{c}(\sigma_{K})$. The morphism $\alpha$ defines an edge to the initial vertex of $Y(d^{c}(\sigma_{K}))$. The extension of this edge yields an $n$-simplex, which we may describe as the composition $\beta\circ\alpha$. We now set $C_{X,Y}\coloneqq\beta\circ\alpha$. Since by design $f$ is cartesian this construction is functorial and yields the desired functor $F\colon\sigma\mapsto C_{X,Y}$. The composition of $F$ with ${\mathcal{D}}$ corresponds to maps $f_{k!}\colon X(k)\rightarrow Y(k)$ compatible with the $X_{kj}^{*}$ resp. $Y_{kj}^{*}$ for all $k\rightarrow j\in K$ (also in higher dimensions by construction). In particular, taking colimits inside $\Pr^{L}$ yields a colimit preserving functor $f_{!}\colon{\mathcal{D}}_{K}(X)\rightarrow{\mathcal{D}}_{K}(Y)$, as desired.
The rest follows from Lemma 2.19 similar to the proof of Proposition 2.18.
@@
Note that it is clear from the proof of Proposition 2.29 that if ${\mathcal{D}}^{!}_{|{\mathcal{C}}_{E}}$ commutes with colimits that also $f^{!}\coloneqq{\mathcal{D}}_{K}^{!}(f)$ commutes with colimits for all $f\in E_{K}$ as ${\mathcal{D}}_{K}(X)$ is compactly generated for all $X\in\operatorname{Fun}(K,{\mathcal{C}})$ by Lemma 2.19.
Proposition 2.31.
Let ${\mathcal{I}}\simeq\mathbb{N}_{0}$, the $\infty$-category associated to the poset $\mathbb{N}_{0}$ seen as a $1$-category. Let $(x,X)$ be a monoidal $-{\mathcal{I}}$-algebraic stack, then
$\textup{DM}(x,X)\simeq\mathop{\textup{colim}}_{i\in{\mathcal{I}},x^{!}}\textup{DM}^{!}(X_{i}).$
Proof.
By descent, we see that for each $i\in{\mathcal{I}}$ we have $\tilde{p}_{i}\colon\textup{DM}^{!}(X_{i})\simeq\textup{DM}^{*}(X_{i})$ as $X_{i}$ admits a smooth covering by a scheme. We will construct equivalences
$p_{i}\colon\textup{DM}^{!}(X_{i})\simeq\textup{DM}^{*}(X_{i})$
out of $\tilde{p}_{i}$ that are compatible with the presentation as a colimit. This is easily achieved by twisting with the Thom-motives associated to the smooth maps $x_{0i}$. But let us be more precise.
Let us define $p_{i}\coloneqq\tilde{p}_{i}\otimes\textup{Th}(\Omega_{x_{0i}})1_{X_{i}}$, where Th denotes the Thom-motive [^1] Let $p\colon V\rightarrow X_{i}$ be a vector bundle with zero section $s$, then $\textup{Th}(V)\coloneqq p_{\sharp}s_{*}$. associated to $\mathbb{V}(\Omega_{x_{0i}})\rightarrow X_{i}$. Then $p_{i}$ is still an equivalence of $\textup{DM}^{!}(X_{i})$ and $\textup{DM}^{*}(X_{i})$. By compatibility with the purity equivalence with composition (cf. [CD19, Rem. 2.4.52]), we claim that this is also compatible with the presentation of $\textup{DM}(x,X)$ as a colimit.
Indeed, the functor $\textup{Spine}[\mathbb{N}_{0}]\rightarrow\mathbb{N}_{0}$ is inner anodyne [Lur21, 03HK], where $\textup{Spine}[\mathbb{N}_{0}]$ is the simplicial subset of $\mathbb{N}_{0}$ of finite vertices that are joined by edges (or more informally, we forget about all $n\geq 2$ simplices in $\mathbb{N}_{0}$). Moreover, by loc. cit., we see that the restriction $\operatorname{Fun}(\mathbb{N}_{0},\textup{DGCat}_{\textup{cont}})\rightarrow\operatorname{Fun}(\textup{Spine}[\mathbb{N}_{0}],\textup{DGCat}_{\textup{cont}})$ is a trivial Kan fibration. Thus, we can lift the map that we constructed to a map of diagrams indexed by $\mathbb{N}_{0}$ proving the claim. ∎
@@
Under the assumption of Proposition 2.31, we can use Proposition 2.18 and similar methods as in the proof to conclude that ${\mathcal{D}}_{{\mathcal{I}}}$ is part of a six functor formalism with respect to the class $\tilde{E}_{{\mathcal{I}}}$, where $\tilde{E}_{{\mathcal{I}}}$ is the class of all morphisms $f$ such that each square is right adjointable and $f_{i}\in E$ for all $i\in{\mathcal{I}}$. To be more precise, we get the existence of $f_{!}$ and $f^{!}$ for morphisms in $\tilde{E}_{{\mathcal{I}}}$ satisfying base change and projection formula.
@@
To generalize Proposition 2.31 it is enough to know that the level-wise equivalences $p_{i}$ in the proof yield an equivalence of diagrams. While on the homotopy categorical level this is clear, we were not able to write a map down for higher simplices.
Proof of Theorem 5.
By Ayoub’s thesis resp. Cisinski-Deglise (cf. [Ayo07] resp. [CD19]) there is a full motivic six functor formalism on Beilinson motives for finite type $S$-schemes. Further, by Robalo’s thesis [Rob15, §2.4], we have that the $6$-functor formalism exists in the sense of [LZ12]. Note that DM also satisfies étale descent. Thus, the proof of the theorem boils down to Proposition 2.31, Proposition 2.29, Remark 2.32, Lemma 2.22, Lemma 2.20 and Lemma 2.19, similar to the proof of Proposition 2.18. ∎
Remark 2.34.
The proof of Theorem 5 can be generalized to other settings such as ind-schemes. To be more specific, using the result of Richarz and Scholbach on rational motives of ind-schemes [RS20, Thm. 2.4.2], we can replace algebraic stacks of finite type, by ind-schemes of ind-finite type and modify DM as follows.
If $(x,X)$ is a $-{\mathcal{I}}$-system of ind-schemes of ind-finite type (let us call such systems $-{\mathcal{I}}$-schemes), then we define $\textup{DM}(x,X)\coloneqq\mathop{\textup{colim}}_{i\in{\mathcal{I}},x^{!}}\textup{DM}^{!}(X_{i})$, note that we use !!!-pullbacks as transition maps.
From now on assume that the transition maps of $-{\mathcal{I}}$-schemes are schematic smooth. Then $\textup{DM}(x,X)$ admits a $\otimes$-product and internal $\operatorname{Hom}$. The existence of the $\otimes$-unit fails in general (cf. loc. cit.). Let $f\colon(x,X)\rightarrow(y,Y)$ be morphism of $-{\mathcal{I}}$-schemes (we define this similarly to Definition 2.2). Assume each $f_{i}$ satisfies the assumptions of [RS20, Thm 2.4.2] for the existence of $f_{i}^{*}$. Then there exists an adjunction $f^{*}\dashv f_{*}$. In general there exists an adjunction $f_{!}\dashv f^{!}$, where we use that $(x,X)$ and $(y,Y)$ are required to be schematic smooth (in this way the right adjoint of $x^{!}$ is up to twist and shift given by $x_{*}$). All of these functors satisfy the properties of Theorem 5 using the same proof.
3 Rational motivic cohomology on pro-algebraic stacks
Before we continue with the definition of motivic cohomology, we want to insert some consequences of Theorem 5 and show that $\textup{DM}(x,X)\simeq\textup{DM}(X)$ for classical monoidal $-{\mathcal{I}}$-algebraic stacks.
Proposition 3.1.
Let $(x,X)$ be a classical monoidal $-{\mathcal{I}}$-algebraic stack. Then we have an equivalence
$\textup{DM}(X)\simeq\textup{DM}(x,X)$
of DG-categories.
Proof.
By assumption $X_{0}$ admits a smooth cover by affine finite type $S$-schemes $U_{0,i}=\operatorname{Spec}(A_{i})$, for some set $I$ and $i\in{\mathcal{I}}$. As $X_{i}\rightarrow X_{0}$ is affine, the pullback $U_{n,i}\coloneqq X_{i}\times_{X_{0}}U_{0,i}$ is also a smooth cover by affines. The same holds after passage to limits, i.e. for $U_{\infty,i}\coloneqq X\times_{X_{0}}U_{0,i}$ the projection $\coprod_{i\in{\mathcal{I}}}U_{\infty,i}\rightarrow X$ is an effective epimorphism. In particular, as DM satisfies $h$-descent (cf. [RS20, Thm. 2.2.16]), we have $\lim_{\Delta}\textup{DM}(\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{\infty,i}/X))\simeq\textup{DM}(X)$. By construction, each of the $U_{\infty,i}$ is equivalent to $\lim_{i\in{\mathcal{I}}}U_{n,i}$. As the $U_{n,i}$ are affine of finite type over $S$, we have
$\textup{DM}(U_{\infty,i})\simeq\textup{DM}(\lim_{i\in{\mathcal{I}}}U_{n,i})\simeq\mathop{\textup{colim}}_{i\in{\mathcal{I}},x^{*}}\textup{DM}(U_{n,i})\simeq\lim_{i\in{\mathcal{I}}^{\operatorname{op}},x_{*}}\textup{DM}(U_{n,i})$
(cf. Lemma 2.20). This yields the equivalence
$\textup{DM}(\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{\infty,i}/X))\simeq\lim_{i\in{\mathcal{I}},x_{*}}\textup{DM}(\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{n,i}/X_{i})).$
Putting all of this together, we claim that this yields
$\textup{DM}(X)\simeq\lim_{\Delta}\textup{DM}(\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{\infty,i}/X))$ $\simeq\lim_{i\in{\mathcal{I}}^{\operatorname{op}},x_{*}}\lim_{\Delta}\textup{DM}(\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{n,i}/X_{i}))$
as desired.
$\textup{DM}(\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{n,i}/X_{i}))\rightarrow\textup{DM}(\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{n-1,i}/X_{j})).$
But this follows immediately from smooth base change, as
$\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{n,i}/X_{i})\simeq\check{C}_{\bullet}(\coprod_{i\in{\mathcal{I}}}U_{n-1,i}/X_{j})\times_{X_{j}}X_{i}.$
Remark 3.2 (Underling motive of pro-algebraic stacks).
Let us show how to compute the underlying motive in $\textup{DM}(S)$ of a monoidal strict $-{\mathcal{I}}$-algebraic stack.
Let $(x,X)$ be a monoidal strict $-{\mathcal{I}}$-algebraic stack. We have seen in Lemma 2.20 that there is a map $p_{0}\colon\textup{DM}(x,X)\rightarrow\textup{DM}(X_{0})$ such that the identity on $\textup{DM}(X_{0})$ factors as
$\textup{DM}(X_{0})\xrightarrow{\textup{ins}_{0}}\textup{DM}(x,X)\xrightarrow{p_{0}}\textup{DM}(X_{0}),$
where $\textup{ins}_{0}$ denotes the natural inclusion.
Assume that ${\mathcal{I}}\simeq\mathbb{N}_{0}$, then by Theorem 5 the two morphisms above correspond to the adjunction
${c_{0*}\colon\textup{DM}(x,X)}$${\textup{DM}(X_{0})\colon c_{0}^{*}}$
where $c_{0}\colon(x,X)\rightarrow X_{0}$ denotes the projection, where we see $X_{0}$ as a constant $-{\mathcal{I}}$-algebraic stack. In particular, since $(x,X)$ is monoidal, both $c_{0}^{!}$ and $c_{0}^{*}$ are fully faithful.
Let us denote by $f\colon X\rightarrow S$ and $f_{0}\colon X_{0}\rightarrow S$ the structure maps and let $M\in\textup{DM}(S)$. The above shows that $f_{!}f^{!}M\simeq f_{0!}f_{0}^{!}M$.
Remark 3.3 (Base change for pro-algebraic stacks).
Let $f\colon(x,X)\rightarrow(y,Y)$ be a smooth morphism monoidal $-{\mathcal{I}}$-algebraic stacks. Assume that $f$ is either adjointable and ${\mathcal{I}}\simeq\mathbb{N}_{0}$ or that $f$ is cartesian and ${\mathcal{I}}$ has an initial object 00.
Our computations have shown that we have $f_{*}\textup{ins}_{0}^{X}\simeq\textup{ins}_{0}^{Y}f_{0*}$, where $\textup{ins}_{0}^{X}$ denotes the natural functor $\textup{DM}(X_{0})\rightarrow\textup{DM}(x,X)$ (similarly for $Y$). This can be see as a form of non-finite type base change equivalence of the square
${(x,X)}$${X_{0}}$${(y,Y)}$${Y_{0},}$
which is cartesian if $f$ is cartesian.
For example this yields a *-base change formula for pullback along the generic point of an integral scheme (cf. Example 3.13 below). This is not true in rational motives, if the inclusion of the generic point may not be of finite type.
Remark 3.4.
@@ 3 Rational motivic cohomology on pro-algebraic stacks
Let $f\colon(x,X)\rightarrow(y,Y)$ be a smooth morphism monoidal $-{\mathcal{I}}$-algebraic stacks. For $M\in\textup{DM}(y,Y)$, we can compute the "motivic global sections with values in $M$" directly. To be more precise, let us note that since $f_{*}f^{*}$ is colimit preserving, its values are determined by its restriction to compacts. As any compact in $\textup{DM}(y,Y)$ comes from a compact in some $\textup{DM}(Y_{i})$ (cf. Lemma 2.19), we see that
@@
In particular, if for example $(y,Y)\cong S$, i.e. the constant diagram associated to $S$, and $(x,X)$ is strict, we have
@@
(cf. Remark 3.2).
3.1 Motivic cohomology

### v4

@@ 3.2 Six functor formalism
First, let us note that the morphism $f$ corresponds to a diagram ${\mathcal{I}}^{\operatorname{op}}\times\Delta^{1,\operatorname{op}}\rightarrow{\mathcal{C}}$. The idea of the proof is to take $f_{!}\coloneqq\mathop{\textup{colim}}_{{\mathcal{I}}}f_{!\bullet}$. However, as $\mathcal{I}$} is arbitrary, we do not immediately obtain a a diagram $f_{!\bullet}\colon{\mathcal{I}}\times\Delta^{1,\operatorname{op}}\rightarrow\textup{DGCat}_{\textup{cont}}$ as in the proof of the $(\ast)$ case. Instead, we will make use of the fact that $f$ is cartesian and we will construct a diagram ${\mathcal{I}}\times\Delta^{1,\operatorname{op}}\rightarrow\textup{Corr}({{\mathcal{C}}})_{E}$. Taking the colimit of this diagram will then yield $f_{!}$. To make this process homotopy coherent, we will use the description [^3] We were made aware of this alternative description by Chirantan Chowdhury. of $\textup{Corr}({{\mathcal{C}}})_{E}$ in terms of bisimplicial sets.
Let us consider the marked simplicial set $({\mathcal{C}},E,\textup{ALL})$, where $E$ is as before and ALL denotes the class of all edges in $\mathcal{C}$}. By the assignment $([n_{1}],[n_{2}])\mapsto\operatorname{Hom}_{\mathrm{Set}_{\Delta}}(\Delta^{n_{1}}\times\Delta^{n_{2}},{\mathcal{C}})$ we obtain a bisimplicial set $\delta_{*}{\mathcal{C}}$. We denote by ${\mathcal{C}}^{\mathrm{cart}}_{E}$ the bisimplicial subset such that the vertical edges are contained in $E$ and each square is cartesian (for a rigorous construction see [Cho25]). By a similar construction, we obtain a bisimplicial set $({\mathcal{I}}^{\operatorname{op}}\times\Delta^{1,\operatorname{op}})^{\textup{ALL}}_{\textup{ALL}}$ associated to the marked simplicial set $({\mathcal{I}}^{\operatorname{op}}\times\Delta^{1,\operatorname{op}},\textup{ALL},\textup{ALL})$. This process is functorial in marked simplicial sets with marked squares, so we obtain a map $\alpha\colon({\mathcal{I}}^{\operatorname{op}}\times\Delta^{1,\operatorname{op}})^{\textup{ALL}}_{\textup{ALL}}\rightarrow{\mathcal{C}}^{\mathrm{cart}}_{E}$. We take opposites of the horizontal arrows, which we see as a functor $\textup{op}^{2}$ and apply this to $\alpha$. Finally, we let $K\mathrm{pt}^{n}$ denote the bisimplicial subset of $\Delta^{(n,n)}$ spanned by edges $(k,l)$ for $0\leq k\leq l\leq n$. Then we define the simplicial set by
$\delta^{*}_{2\nabla}\operatorname{op}^{2}{\mathcal{C}}^{\mathrm{cart}}_{E}\colon n\mapsto\operatorname{Hom}_{\mathrm{Set}_{2\Delta}}(K\mathrm{pt}^{n},\operatorname{op}^{2}{\mathcal{C}}^{\mathrm{cart}}_{E}).$
By functoriality of the construction, we obtain a diagram
$\tilde{\alpha}\colon\delta^{*}_{2\nabla}\operatorname{op}^{2}({\mathcal{I}}^{\operatorname{op}}\times\Delta^{1,\operatorname{op}})^{\textup{ALL}}_{\textup{ALL}}\rightarrow\delta_{2\nabla}^{*}\operatorname{op}^{2}{\mathcal{C}}^{\mathrm{cart}}_{E}.$
It is not hard to see from the construction that $\delta^{*}_{2\nabla}\operatorname{op}^{2}{\mathcal{C}}^{\mathrm{cart}}_{E}\cong\textup{Corr}({{\mathcal{C}}})_{E}$ as simplicial sets. The simplicial set $\delta^{*}_{2\nabla}\operatorname{op}^{2}({\mathcal{I}}^{\operatorname{op}}\times\Delta^{1,\operatorname{op}})^{\textup{ALL}}_{\textup{ALL}}$ is very explicit and we can construct a map of simplicial sets ${\mathcal{I}}\times\Delta^{1,\operatorname{op}}\rightarrow\delta^{*}_{2\nabla}\operatorname{op}^{2}({\mathcal{I}}^{\operatorname{op}}\times\Delta^{1,\operatorname{op}})^{\textup{ALL}}_{\textup{ALL}}$ in the following way.
Note that we have a map
$\operatorname{Hom}_{\mathrm{Set}_{\Delta}}(\Delta^{n}\times\Delta^{n,\operatorname{op}},{\mathcal{I}}^{\operatorname{op}}\times\Delta^{1,\operatorname{op}})\rightarrow(\delta^{*}_{2\nabla}\operatorname{op}^{2}({\mathcal{I}}^{\operatorname{op}}\times\Delta^{1,\operatorname{op}})^{\textup{ALL}}_{\textup{ALL}})_{n}.$
Let $\Delta^{n}\to{\mathcal{I}}\times\Delta^{1,\operatorname{op}}$ be a map of simplicial sets. This map is uniquely determined by projections to $\mathcal{I}$} and $\Delta^{1,\operatorname{op}}$. Let us denote these maps by $p_{1}\colon\Delta^{n}\to{\mathcal{I}}$ and $p_{2}\colon\Delta^{n}\to\Delta^{1,\operatorname{op}}$. From these two maps, we get a map $\Delta^{n}\times\Delta^{n,\operatorname{op}}\to{\mathcal{I}}^{\operatorname{op}}\times\Delta^{1,\operatorname{op}}$ of simplicial sets via
$\varphi_{1}$ $\colon\Delta^{n}\times\Delta^{n,\operatorname{op}}\xrightarrow{\mathrm{pr}_{2}}\Delta^{n,\operatorname{op}}\xrightarrow{p_{1}^{\operatorname{op}}}{\mathcal{I}}^{\operatorname{op}},$
$\varphi_{2}$ $\colon\Delta^{n}\times\Delta^{n,\operatorname{op}}\xrightarrow{\mathrm{pr}_{1}}\Delta^{n}\xrightarrow{p_{2}}\Delta^{1,\operatorname{op}},$
where $\mathrm{pr_{i}}$ denotes the $i$-th projection. In particular, we thus obtain a morphism of simplicial sets
$\phi\colon{\mathcal{I}}\times\Delta^{1,\operatorname{op}}\rightarrow\delta^{*}_{2\nabla}\operatorname{op}^{2}({\mathcal{I}}^{\operatorname{op}}\times\Delta^{1,\operatorname{op}})^{\textup{ALL}}_{\textup{ALL}}\rightarrow\textup{Corr}({{\mathcal{C}}})_{E}.$
@@
By design, we have $\widetilde{{\mathcal{D}}}(X)\simeq{\mathcal{D}}_{{\mathcal{I}}}(X)$, where we view $X$ as an objet of the left hand side via the map $\operatorname{Fun}({\mathcal{I}},{\mathcal{C}}^{\operatorname{op}})^{\operatorname{op}}\to\operatorname{Fun}({\mathcal{I}},\textup{Corr}({{\mathcal{C}}})_{E})^{\operatorname{op}}$. In particular, by applying $\widetilde{{\mathcal{D}}}$ to $\phi$, which we view as an edge in $\operatorname{Fun}({\mathcal{I}},\textup{Corr}({{\mathcal{C}}})_{E})^{\operatorname{op}}$, we obtain a colimit preserving map that preserves compact objects
$f_{!}\coloneqq\widetilde{{\mathcal{D}}}(\phi)\colon{\mathcal{D}}_{{\mathcal{I}}}(X)\rightarrow{\mathcal{D}}_{{\mathcal{I}}}(Y).$
The remaining assertions follow by Lemma 3.8 similar to the proof in the $(\dagger)$ case. ∎
@@
In the proof of Proposition 3.17, we can apply the same method to obtain a diagram ${\mathcal{I}}\times\Delta^{1}\to\textup{Corr}({{\mathcal{C}}})_{E}$ by working with the projection $\Delta^{n}\times\Delta^{n,\operatorname{op}}\to\Delta^{n,\operatorname{op}}.$ The colimit along this map then yields the pullback functor $f^{*}$.
@@
As is evident from the proof of Proposition 3.17, if ${\mathcal{D}}^{!}_{|{\mathcal{C}}_{E}}$ commutes with colimits, then so does
$f^{!}\coloneqq{\mathcal{D}}_{{\mathcal{I}}}^{!}(f)$
for all $f\in E_{{\mathcal{I}}}$. This follows from the fact that ${\mathcal{D}}_{{\mathcal{I}}}(X)$ is compactly generated for all $X\in\operatorname{Fun}({\mathcal{I}},{\mathcal{C}})$, by Lemma 3.8.
Proof of Theorem 2.9.
This is now just a consequence of Proposition 3.16, Proposition 3.17 and Lemma 3.11. Note that if $f$ is a proper cartesian morphism of pro-$\mathcal{I}$}-algebraic stacks, then the construction of $f_{!}$ in Proposition 3.16 shows that $f_{!}\simeq f_{*}$, using that on compacts they agree by Lemma 3.8. ∎
@@
Let $f\colon(x,X)\rightarrow(y,Y)$ be a smooth morphism pro-$\mathcal{I}$}-algebraic stacks. For $M\in\textup{DM}(y,Y)$, we can compute the "motivic global sections with values in MM" directly. To be more precise, let us note that since $f_{*}f^{*}$ is colimit preserving, its values are determined by its restriction to compacts. As any compact in $\textup{DM}(y,Y)$ comes from a compact in some $\textup{DM}(Y_{i})$ (cf. Lemma 3.8), we see that
@@
In particular, if for example $(y,Y)\cong S$, and $(x,X)$ is strict, we have
@@
as expected (cf. Remark 2.15).
Remark 3.21 (Non-finite type base change).
Let $f\colon(x,X)\rightarrow(y,Y)$ be a smooth adjointable morphism pro-algebraic stacks. Assume that $f$ is either ${\mathcal{I}}\simeq\mathbb{N}_{0}$ or that $f$ is cartesian and $\mathcal{I}$} has an initial object 0.
Our computations have shown that we have $f_{*}\textup{ins}_{0}^{X}\simeq\textup{ins}_{0}^{Y}f_{0*}$, where $\textup{ins}_{0}^{X}$ denotes the natural functor $\textup{DM}(X_{0})\rightarrow\textup{DM}(x,X)$ (similarly for $Y$). This can be seen as a form of non-finite type base change equivalence of the square
[…]
which is cartesian if $f$ is cartesian. However, note that as longs as $(x,X)\rightarrow X_{0}$ and $(y,Y)\rightarrow Y_{0}$ are not adjointable, we cannot use the base change of Theorem 2.9.
4 Motivic cohomology of pro-algebraic stacks